OpenStudy (anonymous):

The equation of a line with gradient m is y=m(x-2)+3. If m<0 such that the line meets the positive axis at A and y-axis at B,(a)find the coordinates of A and B. (b)the values of m for which the area enclosed by this line and the axes in 16 units.

OpenStudy (anonymous):

I did part (a). A=(2m-3)/m and 0. B=0 and 3-2m. I need help with part (b)

OpenStudy (experimentx):

y = m(x-2) + 3 = mx + (3 - 2m) ------------------ y - mx = (3-2m) y x -- + ------- = 1 (3 - 2m) (3 - 2m)/(-m) --------------------------- A = (2m-3)/m and B=3-2m ------------------------ 1/2*A*B = 16 -------------- find the value of m

OpenStudy (anonymous):

Thank you so much! ^_^

OpenStudy (experimentx):

yw